OpenStudy (anonymous):

Write the following equation in slope-intercept form. 8x + 4y = 16

OpenStudy (studygurl14):

slope-intercept form y= mx + b 8x + 4y = 16 4y = -8x + 16 y = (-8/4)x + 4 y = -2x + 4